OpenStudy (anonymous):

if x=6sin theta, use trimetric substitution to re write sq rt. 36- x^2

OpenStudy (anonymous):

then replace \(x\) by \(6\sin(\theta)\) and see what you get first step is \[\sqrt{36-(6\sin(\theta))^2}=\sqrt{36-36\sin^2(\theta)}\]

OpenStudy (anonymous):

factor the \(36\) out of the whole thing as \[6\sqrt{1-\sin^2(\theta)}\] and then use a famous trig identity to rewrite \(1-\sin^2(\theta)=\cos^2(\theta)\) and you are almost done

OpenStudy (anonymous):

so its 6 cos theta ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanl you

OpenStudy (anonymous):

yw